OpenStudy (rainbow_dash):

On the Moon, a falling object falls just 2.65 feet in the first second after being dropped. Each second it falls 5.3 feet farther than in the previous second. How far would an object fall in the first ten seconds after being dropped?

OpenStudy (barrycarter):

Hint: How far does it fall in the 2nd second? the 3rd second? Or are you looking for a shortcut?

OpenStudy (anonymous):

\(2.65+2.65+5.3+2.65+2\times 5.3+...\)

OpenStudy (rainbow_dash):

I'm just clueless with this arithmetic stuff :/ the simple sequences i'm fine with, but this junk no way

OpenStudy (anonymous):

do you know what \[\sum_{n=1}^{9}n\]is ?

OpenStudy (rainbow_dash):

Dang if i didn't close the last question then yeh. but not off my mind right now no

OpenStudy (anonymous):

ok it is \[\sum_{k=1}^9=\frac{9\times 10}{2}\] general formula is \[\sum_{k=1}^n=\frac{n(n+1)}{2}\]

OpenStudy (anonymous):

you have a bunch of stuff to add, but 10 terms will contain \(2.65\) and \(10\times 2.65=26.5\)

OpenStudy (anonymous):

then the rest will look like \[\sum_{k=1}^95.3k\] which is the same as \[5.3\sum_{k=1}^9k=5.3\times \frac{9\times 10}{2}\]

OpenStudy (rainbow_dash):

so 5.3*45. which = 238.5